Summary Title: EVAs #10 and #11
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Crew 269 experienced another day of beautiful weather on Mars with mild temperatures, sunny skies, and light high-altitude clouds. Green Hab Officer made blueberry pancakes for breakfast with bourbon maple syrup. After a morning planning meeting, CDR, XO, and HSO conducted EVA #10. The first objective of the EVA was to traverse on foot to retrieve the mirror coating experiment samples from their exposure location across Cow Dung Rd. from Marble Ritual and return the samples to the airlock for analysis under microscope in the Science Dome by Engineer. The EVA then proceeded on rovers to Kissing Camel Ridge, then on foot up to Hab Ridge Rd. for the purpose of regolith sample collection and radiation environment monitoring. After EVA #10 crew returned to the Hab, Engineer and Green Hab Officer embarked on EVA #11, traversing on foot to return the mirror coating experiment samples to their exposure location across Cow Dung Rd. from Marble Ritual. The EVA then proceeded on foot to Sagan St. for the purpose of a handheld ham radio range test with the Hab. The handheld ham radios maintained clear communications on the 2 meter band at a further distance from the Hab than was possible using the EVA radios. During the afternoon EVA, HSO conducted in-depth troubleshooting of the Hab toilet (details can be found in the operations report for today), and Technology Officer conducted troubleshooting of the VR/AR headset in preparation for future EVAs. In the early evening, Green Hab Officer harvested microgreens, HSO collected crew health stats, and the crew took turns completing their exercise protocols. For dinner, the crew enjoyed chicken jambalaya with homemade sourdough bread and a microgreen salad. After dinner, the crew met to discuss prioritizing the remaining mission objectives to ensure maximum science return with the time remaining.